OpenStudy (anonymous):

How many pounds of almonds selling for .85 cents a pound must be mixed with peanuts selling for .55 cents a pound to make a mixture of 50 pounds which will sell for .79 cents a pound? How do i do this?

OpenStudy (mertsj):

Let x be the pounds of almonds Then 50-x = pounds of peanuts .85x+.55(50-x)=.79(50)

OpenStudy (mertsj):

Solve that equation.

OpenStudy (anonymous):

Huh?

hero (hero):

lol, I'm not surprised you don't understand. That's not a very good explanation.

OpenStudy (mertsj):

Can you solve the equation I posted or not?

OpenStudy (mertsj):

Would you agree that if there a x pounds of almonds, then there are 50-x pounds of peanuts?

hero (hero):

@Mertsj, sorry, but I don't really think that matters. The fact that she doesn't understand how you got that equation outweighs solving it.

OpenStudy (anonymous):

A: a+p=50 .85a+.55p=.79 B: a+p=.79(50) .85a+.50p=50 C: a+p=50 .85a+55p=.79(50) D: a+p=50 .55a+.85p=.79(50)

OpenStudy (mertsj):

So God has spoken and God can finish explaining to the asker.

hero (hero):

@Pamelaa23, don't post that. Now someone is just going to give you the answer.

OpenStudy (anonymous):

The reason im confused is how they wrote the answers.

OpenStudy (anonymous):

im not looking for just an answer. I have 13 more problems just like this to do, so i need to know how to do it.

hero (hero):

Okay, I'll help you with this one, but I want to see you post another one and try to solve it on your own.

OpenStudy (anonymous):

Ok, thank you.

hero (hero):

Give me a minute. This site is very laggy

OpenStudy (anonymous):

ok

hero (hero):

We're going to have to do this in vyew. I'll send the link to you in a message.

OpenStudy (anonymous):

ok

hero (hero):

Does the link work?

OpenStudy (anonymous):

yeah.

OpenStudy (anonymous):

You still there?

hero (hero):

Yes, I'm here, where are you?

hero (hero):

I completely explained everything and you didn't respond.

OpenStudy (anonymous):

Idk what happen. But i understand what you wrote, so the answer would be c?

hero (hero):

Yes, the answer is c, but I'm not convinced you fully understand because you never responded. I think you muted your volume.

hero (hero):

So now, you're going to have to post another similar problem and demonstrate your understanding.

OpenStudy (anonymous):

Seriously. I told you that my mic is broke.

hero (hero):

And I told you to use the chat to respond. Which you did initially. Why did you stop using it? And why did you leave?

OpenStudy (anonymous):

I DONT KNOW WHAT HAPPEN! My computer didnt respond to that site, so it closed it.

hero (hero):

Well, like I said, you're gonna have to demonstrate your understanding somehow.

OpenStudy (anonymous):

& if you will give me time to type it. im going to! & clearly i dont HAVE to do anything!

OpenStudy (anonymous):

Im going to post it in a new thread.

hero (hero):

I did you a favor by explaining it to you. You don't have to do it, but you should be considerate and do me a favor and let me know that you understood it.

OpenStudy (mertsj):

Oh for heavens sake. Maybe she has something better to do than stroke your ego.

hero (hero):

@Mertsj, nobody is even talking to you anymore. You should have moved on and started helping other students. Unlike you, I actually try to help students understand. All you do is post the solution and leave.